# Entropy-Gated Branching for Efficient Test-Time Reasoning

### v2

@@ 5 Results
Our main results are presented in Table 1, outlining the performance of each model across the five benchmark datasets. EGB demonstrates substantial improvements over standard base inference, achieving an average gain of 22.6% across all models and benchmarks with absolute accuracy gains of 4-10 percentage points (pp) in most settings. The method consistently delivers competitive or superior performance compared to beam search while achieving these results with significantly reduced computational cost and time, as detailed in Section 5.2. Our analysis reveals interesting patterns across different model families, scales, and reasoning domains.
@@ 5.1 Main results
The Qwen models establish strong baseline performance, with even the smallest Qwen3-1.7B model substantially outperforming comparable Llama models across all benchmarks. For instance, Qwen3-1.7B achieves 48.78% on CFA Level I compared to Llama-3.2-1B’s 30.44%, and 70.40% on MATH-500 versus 21.80% for Llama. This performance advantage persists consistently across all model scales. However, the Llama family demonstrates exceptional responsiveness to EGB, achieving an average improvement of 35.3% over standard decoding compared to Qwen’s more modest 9.9% gains. The Llama models show particularly dramatic improvements on certain tasks. For example, Llama-3.2-1B-instruct improves by +18.34pp on CFA Level II, while the stronger baseline Qwen3-1.7B only gains +3.41pp on the same task. Remarkably, EGB’s substantial gains enable Llama models to bridge much of the performance gap with their Qwen counterparts, often matching or approaching Qwen’s performance despite starting from significantly weaker foundations.
@@
EGB’s advantages over beam search increase monotonically with model size. Small models (1-1.7B) show substantial relative gains from EGB (+31.0% over baseline) but remain slightly behind beam search (-1.2pp average). Medium models (3-4B) achieve balanced performance with EGB winning 60% of comparisons against beam search (+2.9% relative advantage). Large models (8B) demonstrate EGB’s strongest performance profile, with a 70% win rate and substantial improvements over beam search (+7.3% relative advantage). This scaling pattern suggests that larger models don’t reduce exploration headroom but rather enhance the quality and effectiveness of exploration itself. The superior performance of larger models with EGB indicates that increased reasoning capability amplifies the benefits of strategic exploration, as these models can both generate more sophisticated alternative approaches and more accurately evaluate their relative merits. While small models benefit from any additional reasoning paths, larger models can more effectively generate diverse, high-quality alternatives and better distinguish between promising and unpromising directions.
Trends in Benchmark Difficulty
EGB shows the greatest impact on tasks where models already have the necessary knowledge but struggle with consistent execution. On benchmarks like GSM8K and the CFA exams, where models show solid baseline competence, EGB provides substantial gains by helping correct reasoning errors and select among viable solution approaches. Even on highly challenging benchmarks like AIME, where baseline scores rarely exceed 22%, EGB drives considerable improvements, pushing models closer to their performance limits. While exploration is naturally constrained when problems go beyond a model’s knowledge, EGB reliably helps models reach their full potential by testing a wide range of reasoning strategies. This allows them to identify and apply their best possible solutions, even at the edge of their capabilities.
The difficulty-correlated performance pattern illuminates EGB’s primary mechanism as systematic optimization of model capabilities rather than knowledge augmentation. On moderate-difficulty tasks, models typically understand problem requirements and possess appropriate solution strategies but may fail due to computational errors or suboptimal path selection. EGB’s diverse reasoning paths provide multiple opportunities to avoid these execution failures. On the most challenging problems, EGB ensures models leverage every available reasoning strategy, maximizing their performance within existing capability constraints.
Figure 3: Average run-time comparison of various inference methods on five benchmarks
@@ 5.2 Computational Efficiency Analysis
Beyond accuracy improvements, EGB delivers substantial computational efficiency advantages over beam search, achieving 31-75% speedup across all evaluated benchmarks, as shown in Figure 3, while maintaining superior or competitive performance. This efficiency gain represents a significant practical advantage that enhances EGB’s value proposition for real-world deployment scenarios. Unlike traditional accuracy-efficiency trade-offs, EGB achieves both improved performance and reduced computational cost simultaneously. For financial reasoning tasks, EGB delivers 33-40% speedups alongside 2-4 percentage point accuracy improvements. Even on challenging mathematical benchmarks where accuracy gains are modest, the substantial time savings provide clear value for time-sensitive deployments.
@@
5.3 Ablation Studies
@@
We explore how the entropy threshold affects model performance on different scales. By analyzing the effect of altering the entropy threshold $H_{\text{thresh}}$ from 1.0 to 4.0 on a subset of 100 questions from MATH500 dataset, Figure 5 shows that $H_{\text{thresh}}=2$ is, on average, the best setting for the Llama models, where performance remains relatively stable within the range of 1.5-3.5. This optimal range highlights that EGB’s effectiveness depends critically on maintaining sufficient branching opportunities during reasoning while avoiding excessive computational overhead. At very low thresholds ($H_{\text{thresh}}<1.5$), the model branches too frequently, often at decision points where the uncertainty is minimal and exploration provides little benefit. This aggressive branching not only increases computational costs but can also introduce noise by exploring fundamentally similar reasoning paths that differ only in minor details or presentation.
@@
We analyze beam size from KK=2 to K=16, revealing that larger beam sizes consistently improve performance, with the most substantial gains occurring between KK=2 and KK=8 across all model scales. The improvement is particularly pronounced for larger models, with Llama-3.1-8B showing dramatic gains from 36.2% to 48.8% accuracy as beam size increases. This pattern confirms that maintaining multiple active reasoning branches helps preserve solution diversity and prevents premature convergence to suboptimal paths. However, the performance plateaus at KK=16, suggesting diminishing returns. This is likely due to a trade-off in PRM-based ranking: as the PRM may favour clusters of similar partial solutions, additional beams may not contribute new information, limiting the benefits of further expansion.
@@
6 Conclusion
We introduced Entropy-Gated Branching (EGB), a selective expansion strategy that dynamically allocates computational resources based on model uncertainty during generation. We addressed the limitation in current test-time compute methods by recognizing that not all generation steps require equal computational investment. By using entropy as a gating mechanism to trigger branching only at high-uncertainty decision points, EGB achieves a rare combination of improved accuracy and reduced computational cost compared to traditional beam search. Our comprehensive evaluation across mathematical reasoning benchmarks demonstrates that EGB consistently delivers 31-75% speedup while maintaining superior or competitive performance to traditional beam search. The method proves particularly effective on moderately challenging tasks where models possess the requisite knowledge but struggle with execution consistency, achieving substantial accuracy improvements on financial and mathematical reasoning problems. By focusing compute on where it matters the most, EGB offers a pragmatic path to robust, cost-aware test-time reasoning at scale.
7 Disclaimer
This work was supported in part by the J.P. Morgan Chase Faculty Research Award, which was prepared for informational purposes in part by the Artificial Intelligence Research Group of JPMorgan Chase & Co. and its affiliates (“JPMorgan”) and is not a product of the Research Department of JPMorgan. JPMorgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy, or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product, or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.

### v3

@@ 5 Results
Our main results are presented in Table 1, outlining the performance of each model across the five benchmark datasets. EGB demonstrates substantial improvements over standard base inference, achieving an average gain of 18.4% across all models and benchmarks with absolute accuracy gains of 3-8 percentage points (pp) in most settings. The method consistently delivers competitive or superior performance compared to beam search while achieving these results with significantly reduced computational cost and time, as detailed in Section 5.2. Our analysis reveals interesting patterns across different model families, scales, and reasoning domains.
@@ 5.1 Main results
The Qwen models establish strong baseline performance, with even the smallest Qwen3-1.7B model substantially outperforming comparable Llama models across all benchmarks. For instance, Qwen3-1.7B achieves 48.78% on CFA Level I compared to Llama-3.2-1B’s 30.44%, and 70.40% on MATH-500 versus 21.80% for Llama. This performance advantage persists consistently across all model scales. However, the Llama family demonstrates exceptional responsiveness to EGB, achieving an average improvement of 31.2% over standard decoding compared to Qwen’s more modest 12.8% gains. The Llama models show particularly dramatic improvements on certain tasks. For example, Llama-3.2-1B-instruct improves by +8.52pp on CFA Level II, while the stronger baseline Qwen3-1.7B gains +11.93pp on the same task. Remarkably, EGB’s substantial gains enable Llama models to bridge much of the performance gap with their Qwen counterparts, often matching or approaching Qwen’s performance despite starting from significantly weaker foundations.
@@
EGB’s advantages over beam search increase monotonically with model size. Small models (1-1.7B) show substantial relative gains from EGB (+26.1% over baseline) but exhibit mixed results against beam search (-0.8pp average). Medium models (3-4B) achieve balanced performance with EGB winning 60% of comparisons against beam search (+3.2% relative advantage). Large models (8B) demonstrate EGB’s strongest performance profile, with a 70% win rate and substantial improvements over beam search (+5.8% relative advantage).
@@
This scaling pattern suggests that larger models do not reduce exploration headroom but rather enhance the quality and effectiveness of exploration itself. The superior performance of larger models with EGB indicates that increased reasoning capability amplifies the benefits of strategic exploration, as these models can both generate more sophisticated alternative approaches and more accurately evaluate their relative merits. While small models benefit from any additional reasoning paths, larger models can more effectively generate diverse, high-quality alternatives and better distinguish between promising and unpromising directions.
Baseline Comparisons
EGB demonstrates consistent advantages over existing decoding strategies. Compared to self-consistency, EGB achieves superior or comparable performance on 83% of benchmark-model combinations. Leveraging Qwen family models, EGB consistently outperforms self-consistency by an average of 3.2pp, with particularly notable gains on MATH-500 (+3.17pp for Qwen3-1.7B) and GSM8K (+5.23pp for Qwen3-1.7B). Against SEGBS, EGB shows clear advantages on harder benchmarks, winning 70% of comparisons on CFA Level II. This demonstrates the value of leveraging external feedback from process reward models over model self-evaluation, particularly for smaller models that may struggle to reliably assess their own reasoning quality. The performance gap becomes more pronounced on complex tasks where accurate self-feedback requires sophisticated capabilities that smaller models often lack. The comparison with beam search reveals EGB’s efficiency advantage: while achieving comparable or superior accuracy on 67% of tasks, EGB does so with substantially lower computational overhead. EGB’s superior performance stems from its selective expansion strategy. Unlike beam search, which expands at every time step and risks selecting low-quality or similar beams through PRM that lead to incorrect trajectories, EGB strategically targets high-entropy regions to diversify reasoning paths more effectively. These improvements demonstrate that EGB’s entropy-guided exploration effectively captures reasoning uncertainty in ways that complementary approaches cannot fully replicate.
Figure 3: (Left) Budget–accuracy scaling across inference methods. Budget: for Self-Consistency, the budget is the total number of sampled solutions; for beam-style methods (SEGBS, Beam Search, and EGB), the budget is the product of beam width and the number of expansions. (Right) Average runtime per benchmark for EGB and Beam Search, as it proves to be the most competitive baseline.
Computational Budget Analysis
To assess the efficiency and scalability of different decoding methods, we evaluate performance across varying computational budgets ranging from 2 to 32. As illustrated in Figure 3, EGB demonstrates superior scaling characteristics compared to alternative approaches. At low budgets (2-4), EGB achieves competitive performance with beam search and SEGBS, trailing beam search by only 0.8pp at budget=2 but surpassing it by 3.0pp at budget=4. This early crossover point highlights EGB’s efficient utilization of limited computational resources. The advantage becomes more pronounced at higher budgets: at budget=32, EGB reaches 85.71% accuracy, outperforming beam search by 3.02pp, SEGBS by 5.51pp, and self-consistency by 6.47pp. Notably, while all methods show performance saturation at higher budgets, EGB maintains a steeper improvement trajectory, suggesting more effective allocation of additional computation. Self-consistency exhibits the most gradual scaling curve, gaining only 12.64pp from budget=2 to budget=32, compared to EGB’s 15.91pp improvement over the same range. This analysis demonstrates that EGB not only achieves superior final performance but does so with better computational efficiency, making it particularly valuable in resource-constrained settings where maximizing accuracy per unit of computation is critical.
@@ 5.2 Computational Efficiency Analysis
Beyond accuracy improvements, EGB delivers substantial computational efficiency advantages over beam search, achieving 31-75% speedup across all evaluated benchmarks, as shown in Figure 3, while maintaining superior or competitive performance. Because Self-Consistency and SEGBS consistently underperform EGB across budgets and benchmarks, we report runtime comparisons only against Beam Search, the strongest alternative. This efficiency gain represents a significant practical advantage that enhances EGB’s value proposition for real-world deployment scenarios. Unlike traditional accuracy-efficiency trade-offs, EGB achieves both improved performance and reduced computational cost simultaneously. For financial reasoning tasks, EGB delivers 33-40% speedups alongside 2-4pp accuracy improvements. Even on challenging mathematical benchmarks where accuracy gains are modest, the substantial time savings provide clear value for time-sensitive deployments.
@@
6 Conclusion
We introduced Entropy-Gated Branching (EGB), a selective expansion strategy that dynamically allocates computational resources based on model uncertainty during generation. We addressed the limitation in current test-time compute methods by recognizing that not all generation steps require equal computational investment. By using entropy as a gating mechanism to trigger branching only at high-uncertainty decision points, EGB achieves a combination of improved accuracy and reduced computational cost. Our comprehensive evaluation across mathematical reasoning benchmarks demonstrates that EGB consistently delivers substantial speedup while maintaining superior performance over other test-time search algorithms. The method proves particularly effective on moderately challenging tasks where models possess the requisite knowledge but struggle with reliable reasoning paths, achieving substantial accuracy improvements on financial and mathematical reasoning problems. By focusing compute on where it matters the most, EGB offers a pragmatic path to robust, cost-aware test-time reasoning at scale.
Limitations
While EGB demonstrates substantial improvements across multiple benchmarks, several considerations warrant acknowledgment. First, the method’s effectiveness depends on the quality of the external process reward model used for branch evaluation. Variations in PRM quality may influence results in different deployment contexts. Second, the entropy threshold requires model-specific tuning, with optimal values differing between general mathematical problems and complex financial reasoning tasks. Automated threshold selection mechanisms could enhance generalizability. Finally, our evaluation focuses on mathematical and financial reasoning; further investigation is needed to assess EGB’s effectiveness on other reasoning domains, such as commonsense reasoning or open-ended problem-solving.
Appendix A Appendix
A.1 Algorithm
We present the EGB algorithm, which shows how our method dynamically allocates computational resources by branching only when the model exhibits high uncertainty, leading to more efficient test-time computation compared to traditional beam search. We also show an example figure of how the PRM model selects the top-performing beams in Figure 6.
Algorithm 1 Entropy-Gated Branching (EGB)
1:Model $M$, Problem $x$, Beam size $K$, Beam width $W$, Entropy threshold $\tau$, Process Reward Model $f_{PRM}$, Maximum steps $T$
2:Best scoring solution
3:Initialize beam set $\mathcal{B}_{0}=\{x\}$ with single beam containing problem $x$
[…]
5:while $t\leq T$ and beams not terminated do
6:  $\mathcal{P}_{t}\leftarrow\emptyset$ $\triangleright$ Initialize candidate pool
7:  for each beam $b\in\mathcal{B}_{t-1}$ do
8:   Compute entropy: $H_{t}^{(b)}=-\sum_{i=1}^{V}p_{i,t}^{(b)}\log_{2}p_{i,t}^{(b)}$
9:   if $H_{t}^{(b)}\leq\tau$ then $\triangleright$ Certain beam
10:     Generate single candidate: $c_{t,1}^{(b)}\sim M(\cdot|b)$
11:     $\mathcal{P}_{t}\leftarrow\mathcal{P}_{t}\cup\{c_{t,1}^{(b)}\}$
[…]
13:     for $j=1$ to $W$ do
14:      Generate candidate: $c_{t,j}^{(b)}\sim M(\cdot|b,\tau_{temp})$ $\triangleright$ With temperature scaling
15:      $\mathcal{P}_{t}\leftarrow\mathcal{P}_{t}\cup\{c_{t,j}^{(b)}\}$
16:     end for
17:   end if
18:  end for
19:  Remove duplicates from $\mathcal{P}_{t}$ using string matching
20:  for each candidate $c\in\mathcal{P}_{t}$ do
21:   Compute PRM score: $s_{PRM}(c)=f_{PRM}(\text{context},c)$
22:  end for
23:  Sort candidates by PRM scores in descending order
24:  $\mathcal{B}_{t}\leftarrow$ top-$K$ candidates from $\mathcal{P}_{t}$
25:  $t\leftarrow t+1$
26:end while
27:return highest scoring beam from $\mathcal{B}_{t}$
Dataset Details
We evaluate our methods on competition and general mathematical reasoning benchmarks of varying difficulty. First, we include an AIME evaluation set drawn from the AI-MO dataset Art of Problem Solving Community (2025). This corpus contains all 90 problems from the 2022–2024 American Invitational Mathematics Examination (AIME I & II each year) Art of Problem Solving Community (2025). Each AIME test is widely viewed as hard in difficulty between grade-school word problems and full Olympiad-style proofs Mathematical Association of America (2025); Wikipedia contributors (2025).
We also assess performance on the MATH benchmark Hendrycks et al. (2021), which consists of competition-level problems (AMC, AIME, and other sources) spanning a wide range of high-school mathematics topics and difficulty levels; for all experiments, we use 500 held-out test questions. Finally, we report results on GSM8K Cobbe et al. (2021), a widely used grade-school math word-problem benchmark designed to probe multi-step arithmetic reasoning in language models, utilizing 1,319 test questions in our experiments.
In addition to general math benchmarks, we also include financial math datasets to test our method in a domain-specific setting. CFA exams are renowned for their rigorous assessment of financial, quantitative, and analytical reasoning, making them an excellent proxy for evaluating model performance in high-stakes, real-world financial decision-making scenarios. Mistakes in the financial questions can lead to significant losses, reinforcing the need for precise and reliable reasoning. As official CFA exam questions are not publicly accessible, this study utilizes CFA mock exams purchased from AnalystPrep AnalystPrep (2024), covering two levels of the CFA program. The dataset comprises multiple-choice and essay questions, each supplemented with corresponding answers, explanations, grading criteria, and metadata indicating the CFA topic associated with each question.
@@ A.1 Algorithm
LLM-based Evaluation
Final answer correctness was determined using a powerful external language model, Gemini 2.5 Flash, as an impartial judge Gemini Team et al. (2024). This approach circumvents the limitations of rigid evaluation methods like exact string matching or regex-based parsers, which are often brittle and can penalize semantically correct answers due to minor stylistic or formatting variations. For each problem, the LLM judge was provided with the original question, the model’s complete generated response, and the ground-truth solution, and was then prompted to assess whether the final answer was correct. This method ensures a robust and fair assessment by focusing on the semantic correctness of the reasoning outcome rather than superficial syntactic differences, providing a more reliable measure of the model’s true problem-solving capabilities.
A.2 Ablation Studies
@@
We explore how the entropy threshold affects model performance on different scales. By analyzing the effect of altering the entropy threshold $H_{\text{thresh}}$ from 1.0 to 4.0 on a subset of 100 questions from MATH-500 dataset, Figure 5 shows that $H_{\text{thresh}}=2$ is, on average, the best setting for the Llama models, where performance remains relatively stable within the range of 1.5-3.5. This optimal range highlights that EGB’s effectiveness depends critically on maintaining sufficient branching opportunities during reasoning while avoiding excessive computational overhead. At very low thresholds ($H_{\text{thresh}}<1.5$), the model branches too frequently, often at decision points where the uncertainty is minimal and exploration provides little benefit. This aggressive branching not only increases computational costs but can also introduce noise by exploring fundamentally similar reasoning paths that differ only in minor details or presentation.
@@
We analyze beam size from KK=2 to KK=16, revealing that larger beam sizes consistently improve performance, with the most substantial gains occurring between KK=2 and KK=8 across all model scales. The improvement is particularly pronounced for larger models, with Llama-3.1-8B showing dramatic gains from 36.2% to 48.8% accuracy as beam size increases. This pattern confirms that maintaining multiple active reasoning branches helps preserve solution diversity and prevents premature convergence to suboptimal paths. However, the performance plateaus at KK=16, suggesting diminishing returns. This is likely due to a trade-off in PRM-based ranking: as the PRM may favour clusters of similar partial solutions, additional beams may not contribute new information, limiting the benefits of further expansion.
@@
A.3 Reduced Uncertainty
Figure 7 presents the uncertainty distribution of Llama-3.1-8B on a CFA Level II question. The top plot shows the model’s behavior under argmax sampling, which ultimately yields an incorrect answer. In contrast, the bottom plot illustrates entropy-aware branching, resulting in a correct solution. Branching is triggered twice as shown in the horizontal bar at the indicated region. At the first branching point, our method select the same token as argmax, where the entropy and varentropy distribution are the same as argmax. However, at second branching point corresponds to a subsequent decline in both entropy and varentropy, suggesting that the model becomes more confident and stable as it continues reasoning with a different token selection. Meanwhile, in the argmax scenario (top plot), entropy and varentropy remain relatively high and fluctuate at the end of the generation, reflecting continuous uncertainty that ultimately leads to an erroneous conclusion.
Figure 6: One example of PRM response
Figure 7: Entropy and varentropy plots along with sampler states of Llama-3.1-8B-instruct on a CFA II question using argmax sampling (top; incorrect answer) and entropy-aware branching (bottom; correct answer). Branching occurs at the start of the indicated region, resulting in a different reasoning path from argmax being selected. Notably, this causes a downtrend in both entropy and varentropy until the next spike, while the argmax plot remains higher and unstable. Branching points are shown as $\blacksquare$ in the red horizontal bar.